# LLM-in-Sandbox Elicits General Agentic Intelligence

### v2

@@ Appendix D Sandbox Capability Classification
Table 15: Pattern matching rules for detecting sandbox capability usage.
@@ Appendix F Prompt for LLM-in-Sandbox
The prompts used in our experiments are shown in Figure 6 and Figure 7. This represents a minimal baseline prompt that establishes the core sandbox interaction protocol. The prompt can be easily adapted to different use cases by modifying the input/output format specifications (e.g., changing the output file path or format), adding domain-specific instructions, or incorporating additional tools.
@@ 5. Save all outputs to /testbed/output directory
- Use execute_bash to run scripts or commands
- Use str_replace_editor to view, create and edit files
- Use submit to finish once you have completed the task

### v3

@@ Appendix D Computer Capability Classification
Table 15: Pattern matching rules for detecting computer capability usage.
@@ Appendix F Prompt for LLM-in-Sandbox
The prompts used in our experiments are shown in Figure 6 and Figure 7. This represents a minimal baseline prompt that establishes the core environment interaction protocol. The prompt can be easily adapted to different use cases by modifying the input/output format specifications (e.g., changing the output file path or format), adding domain-specific instructions, or incorporating additional tools.
@@ 5. Save all outputs to /testbed/output directory
- Use bash to run scripts or commands
- Use file_editor to view, create and edit files
- Use finish to finish once you have completed the task